OpenStudy (anonymous):

Determine whether the function is a linear transformation: T:P_2 -> P_2, T(a_0+a_1x+a_2x^2) = a_1+2a_2x

razor99 (razor99):

pice of cake :p

razor99 (razor99):

@pipa u there

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

You need to prove that \[T(\alpha \vec{v} + \vec{w})=\alpha T(\vec{v})+T(\vec{w})\]

OpenStudy (anonymous):

okkk thanks

OpenStudy (turingtest):

my understanding is that you need to prove two things\[T(\vec u+\vec v)=T(\vec u)+T(\vec v)\]and\[T(c\vec u)=cT(\vec v)\]

OpenStudy (anonymous):

hmmm but i have another issue with it

OpenStudy (anonymous):

see it says T: P_2 -> P_2 but its not so look at it

OpenStudy (turingtest):

I just took that as a typo...

OpenStudy (anonymous):

whtvr its fine I am gonna ask my tutor

OpenStudy (anonymous):

he just came

OpenStudy (turingtest):

So I am no longer needed?

OpenStudy (anonymous):

well if u have an answer i wld love to hear

OpenStudy (anonymous):

but dont stress urself out over this question @TuringTest

OpenStudy (turingtest):

I'm not so swift with this stuff I'll keep staring at what references I have and let you know if it clicks

OpenStudy (anonymous):

kk Thanks @TuringTest I have the answer already. Thanks seriously. That sounded rude but it wasnt suppossed to. Thanks again

OpenStudy (anonymous):

I don't see any mistake?